OpenStudy (anonymous):

The Ramirez family has three children. The sum of their ages is 11. The product of their ages is 24. How old are the children? PLEASE HELP!

OpenStudy (anonymous):

Ages are 6, 4, and 1 6+4=10 +1=11 6x4=24 x1=24